OpenStudy (anonymous):

a new photocopier can make 72 copies in 2 min. When an older photocopier is operating, the two photocopiers together can make 72 copies in 1.5 min. How long would it take the older photocopier to make 72 copies alone?

OpenStudy (anonymous):

If anyone would be willing to help me, I would appreciate it very much.

OpenStudy (anonymous):

I am thinking that the way to solve this problem is as follows: You have the first rate, 72 copies / 2 minutes (that's the new copier) + the old copier (we don't know the rate for that) = 72 copies / 1.5 minutes (the final rate). So the equation should look like this: \[\frac{ 72 }{ 2 } + x = \frac{ 72 }{ 1.5 }\]

OpenStudy (anonymous):

So if we solve that equation for x, we get...

OpenStudy (anonymous):

84?

OpenStudy (anonymous):

12 copies / minute

OpenStudy (anonymous):

72/2 = 36; 72/1.5 = 48, so it is: 36 + x = 48

OpenStudy (anonymous):

oops - didn't read the question thoroughly - I only got the rate of the old copier.

OpenStudy (anonymous):

so for 72 copies @ 12 copies / minute, it would take 6 minutes.

OpenStudy (anonymous):

Does that make sense to you?

OpenStudy (anonymous):

Yes, I think so. Thank you very much for taking the time to explain it to me. :)

OpenStudy (anonymous):

no worries :)

OpenStudy (anonymous):

Have a Nice day.